OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 5 - 15 + 45 - 135 + ...

OpenStudy (anonymous):

first factor out 5 then it will be much easier

OpenStudy (ankit042):

This is a geometric series with r = -3,a=5 Do you knoe the formula to calculate the sum of Geometric series?

OpenStudy (anonymous):

no lol xD what is it?

OpenStudy (anonymous):

5(1-3+9-27...) these are all powers of 3

OpenStudy (anonymous):

1 is zeroth power 3 is first power 9 is secend power and so on

OpenStudy (ankit042):

You can check this video to learn more about Geometric series http://www.youtube.com/watch?v=pXo0bG4iAyg

OpenStudy (anonymous):

\[\sum_{oo}^{n=0}5*3^n ?\]

OpenStudy (anonymous):

oops switch the n=0 and oo

OpenStudy (anonymous):

yes but there is a tiny problem alternating sign

OpenStudy (anonymous):

if you will expend your summation you will get 5(1+3+9+27...)

OpenStudy (anonymous):

you need odd powers to have -1 in front

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}5*\left( -3 \right)^{n}or \sum_{n=0}^{\infty}5*\left( -1 \right)^{n}3^{n}\]

OpenStudy (anonymous):

yes well done

OpenStudy (anonymous):

OHH awesome thank you :)

OpenStudy (anonymous):

@surjithayer could ya give @le0n a medal too!